OpenStudy (anonymous):

An equation is shown below: 4x^2 -1 / 2y - 9y - 5 / 3 = z - 3y "Which is an equivalent equation solved for y" is what it is asking me... I've been working on this for days and I can officially say I have no idea what I'm doing... help! Please and thank you in advance.

OpenStudy (anonymous):

\[4x ^{2}-\frac{ 1 }{ 2y }-9y-\frac{ 5 }{ 3 }=z-3y\]

OpenStudy (anonymous):

right?

OpenStudy (anonymous):

I attached a picture showing it just to make sure there was no confusion, hope it helps :)

OpenStudy (anonymous):

The letter z will be in red as they look like my 2's, but these are the options given.

OpenStudy (anonymous):

D

OpenStudy (anonymous):

Ohhhhhh~ the only difference between the two is ones negative and one is positive... sweet! Do you think you can give a brief explanation on how you got that answer?